### v2

# A Representation Transformation of Parametric Feynman Integrals

Abstract
A transformation on homogeneous polynomials is proposed, which is further applied to parametric Feynman integrals. The two representations related through this transformation are dual to each other. And it naturally leads to dualities of Landau equations and linear integral relations between the two representations. For integrals with momentum-space correspondences, the dual representation is equivalent to the Baikov representation.
I Introduction
Feynman integrals are a crucial element of perturbative quantum field theory. There are several representations of Feynman integrals, with the most well-known being the Feynman-parameter representation [1, 2, 3, 4]. Variants of the Feynman-parameter representation include the Lee-Pomeransky representation [5, 6] and the version used by the author in refs. [7, 8, 9, 10]. The parametric representation is powerful in the sense that it allows for both direct integration [11, 12, 13, 14, 15] and integral reduction [16, 6, 17, 7, 8, 9, 18, 19, 20], and Landau equations [21, 22, 23] can easily be formulated within this framework (see e.g. ref. [24]). Furthermore, the parametric representation can reveal some hidden mathematical structures of Feynman integrals due to its connections with hypergeometric functions [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35], periods [36, 37, 14], coaction [38, 39, 40, 41, 42], etc.
Another widely used representation is the Baikov representation [43, 44, 45, 46, 47] and its variants [48, 49, 50]. Although it is less suited for direct integration, the Baikov representation has the advantage that implementing cuts is straightforward [48, 51, 52], and it is more compatible with twisted cohomology theory [53, 54, 55, 56]. The Baikov representation has been extensively applied to integral reduction [57, 58, 53, 54, 55, 56, 59, 60], constructing differential equation basis [61, 62, 63], Landau analysis [64, 65, 66, 67], etc..
While it is evident that these two representations should be equivalent, since both are derived from the momentum representation, it was not clear how to directly derive one representation from the other. This problem is addressed in this article. It is demonstrated that these two representations are related by a simple transformation. A transformation on homogeneous polynomials is introduced in section II. The transformation is then applied to parametric Feynman integrals in section III. It is proven that the two representations related through this transformation are dual to each other. The equivalence between the dual representation and the Baikov representation is proven in section IV.
II Polynomial transformation
We consider the following transformation on a homogeneous polynomial $\mathcal{F}(x_{1},x_{2},\dots,x_{n+1})$.
$u_{i}\equiv\frac{1}{\mathcal{F}}\frac{\partial\mathcal{F}}{\partial x_{i}},$ (II.1a)
$\mathcal{G}(u)\equiv\left.\frac{1}{\mathcal{F}}\right|_{x=x(u)}.$ (II.1b)
We assume that the degree of $\mathcal{F}$ is $L+1$. Then obviously
$\sum_{i=1}^{n+1}u_{i}x_{i}=L+1.$ (II.2)
Taking the derivative of both sides with respect to $u_{i}$, we get
$x_{i}+\sum_{j=1}^{n+1}u_{j}\frac{\partial x_{j}}{\partial u_{i}}=0.$
On the other hand,
$\begin{split}\frac{1}{\mathcal{G}}\frac{\partial\mathcal{G}}{\partial u_{i}}=\mathcal{F}\frac{\partial}{\partial u_{i}}\left(\frac{1}{\mathcal{F}}\right)=-\frac{1}{\mathcal{F}}\sum_{j=1}^{n+1}\frac{\partial\mathcal{F}}{\partial x_{j}}\frac{\partial x_{j}}{\partial u_{i}}=-\sum_{j=1}^{n+1}u_{j}\frac{\partial x_{j}}{\partial u_{i}}.\end{split}$
A combination of the above two equations leads to
$x_{i}=\frac{1}{\mathcal{G}}\frac{\partial\mathcal{G}}{\partial u_{i}}.$ (II.3)
Thus we see that the transformation defined by eq. II.1 is reciprocal.
Generally speaking, the transformation defined by eq. II.1a may not be reversible. Even if it is, the obtained function $\mathcal{G}(u)$ may not be a rational function. However, for the $\mathcal{F}$ polynomial of a loop integral with a complete set of propagators, $\mathcal{G}$ is indeed a rational function, as will be clear in the subsequent sections.
III Representation transformation
As was shown in refs. [7, 8, 9], a Feynman integral can be parametrized by scalar integrals of the following form.
$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\ldots,\lambda_{n})\\ =&\frac{\Gamma(-\lambda_{0})}{\prod_{i=m+1}^{n+1}\Gamma(\lambda_{i}+1)}\int\mathrm{d}\Pi_{x}^{(n+1)}\mathcal{F}^{\lambda_{0}}\prod_{i=1}^{m}x_{i}^{-\lambda_{i}-1}\prod_{i=m+1}^{n+1}x_{i}^{\lambda_{i}}~{}.\end{split}$ (III.1)
[…]
We introduce the function [9]
$w_{\lambda}(x)\equiv e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\mathrm{d}v~{}\frac{1}{(v-i0^{+})^{\lambda+1}}e^{ivx}~{}.$ (III.2)
We have
$w_{-1}(x)=$ $2\pi\delta(x)~{},$
$w_{\lambda}(x)=$ $2\pi\theta(x)\frac{x^{\lambda}}{\Gamma(\lambda+1)},~{}\lambda\geq 0~{}.$
And we have
$\frac{1}{x^{\lambda+1}}=e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\frac{\mathrm{d}v}{2\pi}~{}w_{\lambda}(x)e^{ivx}~{}.$ (III.3)
Obviously we can write $\frac{1}{\Gamma(\lambda+1)}\int_{0}^{\infty}\mathrm{d}xx^{\lambda}$ as $\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathrm{d}xw_{\lambda}(x)$. By expressing $w_{\lambda}(x)$ (or $x^{-\lambda-1}$) in terms of $v^{-\lambda-1}$ [or $w_{\lambda}(v)$], it is easy to get
$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\dots,\lambda_{n})\\ =&(2i\pi)^{-n-1}\Gamma(-\lambda_{0})\exp\left(-\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}v_{i}\prod_{i=1}^{m}w_{\lambda_{i}}(v_{i})\prod_{i=m+1}^{n+1}v_{i}^{-\lambda_{i}-1}\int_{-\infty}^{\infty}\mathrm{d}\Pi_{x}^{(n+1)}\mathcal{F}^{\lambda_{0}}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)\\ =&(2i\pi)^{-n-1}\Gamma(-\lambda_{0})\exp\left(-\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)\int_{-\infty}^{\infty}\mathrm{d}\Pi_{v}^{(n+1)}\prod_{i=1}^{m}w_{\lambda_{i}}(v_{i})\prod_{i=m+1}^{n+1}v_{i}^{-\lambda_{i}-1}\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)~{},\end{split}$ (III.4)
where the last step can be proven by inserting an integral $\int\mathrm{d}s\delta(s-\mathcal{E}^{(1)}(v))$, rescaling $x$ and $v$ by $x_{i}\to\frac{x_{i}}{s}$ and $v_{i}\to sv_{i}$, and finally integrating out $s$.
We consider a special class of integrals of which the $\mathcal{F}$ polynomials are complete. Here, by complete, we mean that there are $n$ linearly independent matrices $\Lambda_{i,jk}$ such that
$\sum_{j,k=1}^{n+1}\Lambda_{i,jk}u_{j}x_{k}=0,~{}i=1,2,\dots,n~{}.$ (III.5)
We further assume that
$\Lambda_{i,j(n+1)}=$ $0,~{}j\neq n+1~{},$ (III.6a)
$\sum_{j=1}^{n}\Lambda_{i,jj}=$ $n_{0}\Lambda_{i,(n+1)(n+1)}~{},$ (III.6b)
with $n_{0}$ a constant. As can be seen in appendix B, normal loop integrals do satisfy these conditions. The first condition is a consequence of the fact that $\mathcal{F}$ is linear in $x_{n+1}$. For a complete homogeneous $\mathcal{F}$ polynomial satisfying the conditions in eq. III.6, it can be proven that (see appendix A)
$\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)=C_{\mathcal{F}}v_{n+1}^{-n_{0}-1}\mathcal{G}(v)^{-\lambda_{0}^{\prime}}~{},$ (III.7)
where $\lambda_{0}^{\prime}\equiv\lambda_{0}+\frac{n-n_{0}}{L+1}$ and $C_{\mathcal{F}}$ is a constant. Thus we get
$\begin{split}&I_{x}(\lambda_{0},\lambda_{1},\dots,\lambda_{n})\\ =&(2i\pi)^{-n-1}\Gamma(-\lambda_{0})\exp\left(-\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)C_{\mathcal{F}}\\ &\times\int_{-\infty}^{\infty}\mathrm{d}\Pi_{u}^{(n+1)}\mathcal{G}(u)^{-\lambda_{0}^{\prime}}\prod_{i=1}^{m}w_{\lambda_{i}^{\prime}}(u_{i})\prod_{i=m+1}^{n+1}u_{i}^{-\lambda_{i}^{\prime}-1}\\ \equiv&(2i\pi)^{-n-1}\exp\left(-\frac{i\pi}{2}\sum_{i=1}^{n+1}\lambda_{i}\right)C_{\mathcal{F}}\frac{\Gamma(-\lambda_{0})}{\Gamma(\lambda_{0}^{\prime})}\\ &\times I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})~{},\end{split}$ (III.8)
where $\lambda_{n+1}^{\prime}=\lambda_{n+1}+n_{0}+1$, and $\lambda_{i}^{\prime}=\lambda_{i},~{}i\neq 0,n+1$.
Thus, we have proven that $I_{x}(\lambda_{0},\lambda_{1},\dots,\lambda_{n})$ equals $I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})$ up to some constants, and there is an elegant duality between these two representations, which is summarized in table 1.
$\mathcal{G}=\frac{1}{\mathcal{F}}$ $\mathcal{F}=\frac{1}{\mathcal{G}}$
$u=\frac{1}{\mathcal{F}}\frac{\partial\mathcal{F}}{\partial x}$ $x=\frac{1}{\mathcal{G}}\frac{\partial\mathcal{G}}{\partial u}$
$\frac{1}{u^{\lambda+1}}=e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\frac{\mathrm{d}x}{2\pi}~{}w_{\lambda}(x)e^{iux}$  ${}w_{\lambda}(x)\equiv e^{-\frac{\lambda+1}{2}i\pi}\int_{-\infty}^{\infty}\mathrm{d}u~{}\frac{1}{u^{\lambda+1}}e^{iux}$
Table 1: The duality between the two representations.
IV The Baikov representation
We consider a loop integral with a complete set of propagators
$v_{i}=\sum_{j,k}A_{i,jk}l_{j}\cdot l_{k}+2\sum_{j,k}B_{i,jk}l_{j}\cdot p_{k}+C_{i},\quad i=1,~{}2,\dots,n~{},$ (IV.1)
where $l$ are the loop momenta and $p$ are the external momenta. For simplicity, I assume that all the propagators are linearly independent and the Gram determinant of the external momenta $p$ does not vanish. For future convenience, we collectively denote $l$ and $p$ by $q$. Scalar integrals can be parametrized by integrals of the form [44, 45]
$J(\lambda_{0},\lambda_{1},\ldots,\lambda_{n})=\int\prod_{i=1}^{n}\mathrm{d}v_{n}~{}P^{-\lambda_{0}-\frac{1}{2}(L+E+1)}\prod_{i=1}^{n}v_{i}^{-\lambda_{i}-1}~{},$ (IV.2)
where $E$ is the number of the external momenta $p$. Here we have omitted some irrelevant prefactors for brevity. The polynomial $P(v)$ is proportional to the Gram determinant of the momenta $q$. That is
$P(v)\equiv\frac{\det(q_{i}\cdot q_{j})}{\det(p_{i}\cdot p_{j})}~{}.$ (IV.3)
In this section, we will prove that $J(\lambda_{0},\lambda_{1},\ldots,\lambda_{n})$ is equivalent to the integral $I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})$ in eq. III.8 up to some constants.
Taking $\mathcal{E}^{(1)}=|u_{n+1}|$ and $m=0$ in the integral $I_{u}$, we get
$I_{u}(\lambda_{0}^{\prime},\lambda_{1}^{\prime},\dots,\lambda_{n}^{\prime})$
$=$ $\Gamma(\lambda_{0}^{\prime})\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\left[\mathcal{G}_{+}^{-\lambda_{0}^{\prime}}+\mathcal{G}_{-}^{-\lambda_{0}^{\prime}}e^{i\pi(\lambda_{n}^{\prime}+1)}\right]\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
$=$ $\Gamma(\lambda_{0}^{\prime})\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
×$\times\left[\mathcal{G}_{-}^{-\lambda_{0}^{\prime}}e^{-i\pi(\lambda_{n}^{\prime}+1)}+\mathcal{G}_{-}^{-\lambda_{0}^{\prime}}e^{i\pi(\lambda_{n}^{\prime}+1)}\right]$
$=$ $2\Gamma(\lambda_{0}^{\prime})\cos\left(\pi(\lambda_{n+1}^{\prime}+1)\right)\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\mathcal{G}_{-}^{-\lambda_{0}^{\prime}}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}$
$=$ $\frac{-4i}{\Gamma(1-\lambda_{0}^{\prime})}\cos\left(\pi(\lambda_{n+1}^{\prime}+1)\right)$
×$\times\int_{-\infty}^{\infty}\prod_{i=1}^{n}\mathrm{d}u_{i}~{}\theta\left(-\mathcal{G}_{-}\right)\left(-\mathcal{G}_{-}\right)^{-\lambda_{0}^{\prime}}\prod_{i=1}^{n}u_{i}^{-\lambda_{i}^{\prime}-1}~{},$
where $\mathcal{G}_{+}\equiv\left.\mathcal{G}\right|_{u_{n+1}=1}$ and $\mathcal{G}_{-}\equiv\left.\mathcal{G}\right|_{u_{n+1}=-1}$. Due to the positivity of $|u_{n+1}|$, the integral recieves contributions from both $u_{n+1}=1$ and $u_{n+1}=-1$. The second step is carried out by reversing the path of integration for all $x_{i}$ for the first term in the squared bracket. And in the last step, we have closed the contour around the branch cut. Notice that we use such a convention that $\mathrm{Im}\{\mathcal{F}\}<0$. Thus $\mathrm{Im}\{\mathcal{G}\}>0$ when $\mathrm{Re}\{\mathcal{G}\}<0$. Now the region of integration is constrained by the theta function $\theta\left(-\mathcal{G}_{-}\right)$. Since $n_{0}=n-\frac{1}{2}(L+1)(L+E+1)$ (see appendix B), we have $\lambda_{0}^{\prime}=\lambda_{0}+\frac{1}{2}(L+E+1)$. Comparing the above integral with the Baikov integral in eq. IV.2, we see that these two integrals are equivalent if
$P(v_{1},v_{2},\dots,v_{n})=-\mathcal{G}_{-}=-\mathcal{G}(v_{1},v_{2},\dots,v_{n},-1)~{}.$ (IV.4)
Immediately, we will see that this is indeed true.
We define
$A_{ij}\equiv$ $\sum_{k=1}^{n}A_{k,ij}x_{k}~{},$
$B_{ij}\equiv$ $\sum_{k=1}^{n}B_{k,ij}x_{k}~{},$
$C\equiv$ $\sum_{i=1}^{n}C_{i}x_{i}~{}.$
Then we have [7]
$\mathcal{F}=\sum_{i,j,k,l}p_{k}\cdot p_{l}\tilde{A}_{ij}B_{ik}B_{jl}+(x_{n+1}-C)U~{},$ (IV.5)
where
$U\equiv$ $\det(A_{ij})~{},$
$\tilde{A}_{ij}\equiv$ $\left(A^{-1}\right)_{ij}U~{}.$
Let $a_{ij}$, $b_{ij}$, $\alpha_{ij,k}$, and $\beta_{ij}$ be those defined in ref. [8]. That is, they are solutions to the following equations.
$\sum_{l}a_{il}B_{l,jk}=0,$ (IV.6)
$\sum_{l}b_{il}A_{l,jk}=0,$ (IV.7)
$\sum_{kl}\alpha_{ij,k}a_{kl}A_{l,mn}=\frac{1}{2}\left(\delta^{im}\delta^{jn}+\delta^{in}\delta^{jm}\right),$ (IV.8)
$\sum_{kl}\beta_{ij,k}b_{kl}B_{l,mn}=\delta^{im}\delta^{jn}~{}.$ (IV.9)
Then it can be proven that (see appendix B)
$\begin{split}\mathcal{G}=&u_{n+1}\det\left[\sum_{kl}\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})\right.\\ &\left.+u_{n+1}\sum_{mn}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}\right]~{},\end{split}$ (IV.10)
where $u_{i}$ are those appeared in eq. II.1, $g_{ij}$ is the inverse of the Gram matrix $p_{i}\cdot p_{j}$, and
$B^{\prime}_{ij}(u)\equiv\frac{1}{2u_{n+1}}\sum_{k,l}\beta_{ij,k}b_{kl}\left(u_{l}+C_{l}u_{n+1}\right)~{}.$
On the other hand, $q_{i}\cdot q_{j}$ can be understood as a block matrix. Then we have
$P=$ $\frac{1}{\det(p_{i}\cdot p_{j})}\det\begin{pmatrix}l_{i}\cdot l_{j}&l_{i}\cdot p_{j}\\ l_{i}\cdot p_{j}&p_{i}\cdot p_{j}\end{pmatrix}$
$=$ $\det\left(l_{i}\cdot l_{j}-\sum_{mn}g_{mn}l_{i}\cdot p_{m}l_{j}\cdot p_{n}\right)~{}.$
It is easy to get
$l_{i}\cdot p_{j}=$ $\frac{1}{2}\sum_{k,l}\beta_{ij,k}b_{kl}(v_{l}-C_{l})\equiv-B^{\prime\prime}_{ij}=-\left.B^{\prime}_{ij}(v)\right|_{v_{n+1}=-1}~{},$
$l_{i}\cdot l_{j}=$ $\sum_{k,l}\alpha_{ij,k}a_{kl}(v_{l}-C_{l})~{}.$
Thus
$P=\det\left(\sum_{kl}\alpha_{ij,k}a_{kl}(v_{l}-C_{l})-\sum_{mn}g_{mn}B^{\prime\prime}_{im}B^{\prime\prime}_{jn}\right)~{}.$ (IV.11)
Comparing eq. IV.10 with eq. IV.11, we get eq. IV.4.
V Discussion
A reversible transformation on homogeneous polynomials is proposed. The transformation is further applied to parametric Feynman integrals, since a parametric integral is characterized by a homogeneous polynomial. An exact duality between the two representations related through this transformation is proven, as summarized in table 1. Furthermore, it is proven that the dual representation is equivalent to the Baikov representation.
The transformation and the corresponding duality can be applied to various aspects of Feynman integrals, including Laundau analysis and integral reduction. Specifically, the transformation in II.1 establishes a duality between the first (second) Landau equation in the parametric representation and the second (first) Landau equation in the Baikov representation. Additionally, eq. III.5 establishes a correspondence between a syzygy equation in the Baikov representation [57, 58] and a parametric annihilator in the parametric representation [6, 17].
Acknowledgements.
This work is supported by the Natural Science Foundation of China (NSFC) under the contract No. 11975200, and by Guangdong Major Project of Basic and Applied Basic Research (No. 2020B0301030008).
Appendix A Proof of eq. (III.7)
Combining the homogeneity condition eq. II.2 with those conditions in eqs. III.5 and III.6, we have
$0=$ $\sum_{j,k=1}^{n+1}\Lambda_{i,jk}u_{j}x_{k}$
$=$ $\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda_{i,jk}u_{j}x_{k}+\Lambda_{i,(n+1)(n+1)}u_{n+1}x_{n+1}$
$=$ $(L+1)\Lambda_{i,(n+1)(n+1)}$
$+\sum_{j=1}^{n+1}\sum_{k=1}^{n}\left(\Lambda_{i,jk}-\Lambda_{i,(n+1)(n+1)}\delta_{jk}\right)u_{j}x_{k}$
$\equiv$ $(L+1)\Lambda_{i,(n+1)(n+1)}+\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}u_{j}x_{k}~{}.$
For simplicity, we denote
$I_{0}(f)\equiv\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}f(x)\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)~{},$
with $f(x)$ a polynomial in $x$. Then we have
$-(L+1)\lambda_{0}\Lambda_{i,(n+1)(n+1)}I_{0}(1)$
$=$ $\lambda_{0}\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}I_{0}(u_{j}x_{k})$
$=$ $\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\frac{\partial\mathcal{F}^{\lambda_{0}}}{\partial x_{j}}x_{k}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)$
$=$ $-\sum_{j=1}^{n}\Lambda^{\prime}_{i,jj}I_{0}(1)-i\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}v_{j}I_{0}(x_{k})$
$=$ $(n-n_{0})\Lambda_{i,(n+1)(n+1)}I_{0}(1)-\sum_{j=1}^{n+1}\sum_{k=1}^{n}v_{j}\frac{\partial I_{0}(1)}{\partial v_{k}}~{}.$
That is
$\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}v_{j}\frac{\partial I_{0}(1)}{\partial v_{k}}=$ $\left[n-n_{0}+(L+1)\lambda_{0}\right]$
×$\times\Lambda_{i,(n+1)(n+1)}I_{0}(1)~{}.$
Thus
$\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}v_{j}\frac{\partial}{\partial v_{k}}\left[\mathcal{G}(v)^{\lambda_{0}+\frac{n-n_{0}}{L+1}}I_{0}(1)\right]=0~{}.$
By assumption, the matrix $\Lambda^{\prime}_{ij}(v)\equiv\sum_{k=1}^{n+1}\Lambda^{\prime}_{i,kj}v_{k}$ is invertible. Thus $\mathcal{G}(v)^{\lambda_{0}+\frac{n-n_{0}}{L+1}}I_{0}(1)$ is independent of $v_{i},~{}i\neq n+1$. That is
$I_{0}(1)=C_{\mathcal{F}}v_{n+1}^{-n_{0}-1}\mathcal{G}(v)^{-\lambda_{0}-\frac{n-n_{0}}{L+1}}~{},$ (A.1)
where $C_{\mathcal{F}}$ is a constant. The $v_{n+1}$ dependence is fixed through a simple power counting.
Appendix B Derivation of eq. (IV.10)
By virtue of eq. IV.5, we have (In this section, summations over repeated indices are always implied.)
$\beta_{ij,k}b_{kl}u_{l}\mathcal{F}=$ $\beta_{ij,k}b_{kl}\frac{\partial\mathcal{F}}{\partial x_{l}}$
$=$ $2\tilde{A}_{ik}B_{kl}p_{l}\cdot p_{j}-\beta_{ij,k}b_{kl}C_{l}U$
$=$ $2\left(A^{-1}\right)_{ik}B_{kl}p_{l}\cdot p_{j}u_{n+1}\mathcal{F}-\beta_{ij,k}b_{kl}C_{l}u_{n+1}\mathcal{F}~{}.$
Thus we have
$B_{ij}=A_{ik}g_{jl}B^{\prime}_{kl}~{}.$
Similarly,
$\alpha_{ij,k}a_{kl}u_{l}\mathcal{F}=\alpha_{ij,k}a_{kl}\frac{\partial\mathcal{F}}{\partial x_{l}}$
$=$ $-U\left[p_{m}\cdot p_{n}\left(A^{-1}\right)_{ik}\left(A^{-1}\right)_{jl}B_{km}B_{ln}+\alpha_{ij,k}a_{kl}C_{l}\right]$
$+\tilde{A}_{ij}\left[p_{k}\cdot p_{l}\left(A^{-1}\right)_{mn}B_{mk}B_{nl}+x_{n+1}-C\right]$
$=$ $-\mathcal{F}u_{n+1}\left(g_{mn}B^{\prime}_{im}B^{\prime}_{jn}+\alpha_{ij,k}a_{kl}C_{l}\right)+\left(A^{-1}\right)_{ij}\mathcal{F}~{}.$
That is
$\left(A^{-1}\right)_{ij}=\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})+u_{n+1}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}~{}.$
Taking the determinant of both sides, we get
$\begin{split}\mathcal{G}=&\mathcal{F}^{-1}=u_{n+1}U^{-1}=u_{n+1}\det\left(A^{-1}\right)\\ =&u_{n+1}\det\left[\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})+u_{n+1}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}\right]~{}.\end{split}$ (B.1)
Following some similar algebras, we can also get
$0=$ $A_{ik}\beta_{kj,l}b_{lm}(u_{m}+C_{m}u_{n+1})-2B_{ik}p_{k}\cdot p_{j}u_{n+1}~{},$ (B.2a)
$0=$ $(L+1)\left(A_{ik}\alpha_{kj,l}a_{lm}+\frac{1}{2}B_{ik}\beta_{jk,l}b_{lm}\right)$ (B.2b)
×$\times(u_{m}+C_{m}u_{n+1})-\delta_{ij}u_{k}x_{k}~{}.$ (B.2c)
These equations are of the form of the complete condition eq. III.5, since $A_{ij}$ and $B_{ij}$ are linear in $x$. It is easy to see that they contain $\frac{1}{2}L(L+1)+E=n$ independent equations. $\sum_{j=1}^{n}\Lambda_{i,jj}$ can be obtained from eq. III.5 by replaceing $u_{j}x_{k}$ by $\delta_{jk}$. Thus eq. B.2a gives [Remember the definitions in eq. IV.6.]
$\sum_{j=1}^{n}\Lambda_{i,jj}=\Lambda_{i,(n+1)(n+1)}=0~{}.$
And eq. B.2c gives (Notice that the matrix $A_{ij}$ is of dimension $L\times L$, and the matrix $B_{ij}$ is of dimension $L\times E$.)
$\Lambda_{i,(n+1)(n+1)}=$ $-1~{},$
$\sum_{j=1}^{n}\Lambda_{i,jj}=$ $\frac{1}{2}(L+1)(L+1+E)-n~{}.$
That is,
$n_{0}=n-\frac{1}{2}(L+1)(L+1+E)~{}.$ (B.3)